There are 33 coins that equal $6.05 with only quarters and nickels. How many of each do they have?
Answer by mananth(16949)

Nickels x numbers
Quarters y numbers

1 x + 1 y = 33 .............1
5 x + 25 y = 605 .............2
Eliminate y
multiply (1)by -25
Multiply (2) by 1
-25 x -25 y = -825
5 x 25 y = 605
Add the two equations
-20 x = -220
/ -20
x = 11
plug value of x in (1)
1 x + 1 y = 33
11 + 1 y = 33
1 y = 33 -11
1 y = 22
y = 22
Nickels 11 numbers
Quarters 22 numbers
m.ananth@hotmail.ca